OpenStudy (zeidyt):

if (54)m = 512, what is the value of m? 2 3 5 8

OpenStudy (danjs):

multiply both sides by 1/54

OpenStudy (jocelm):

or divide both sides by 54.

OpenStudy (zeidyt):

what do i multiply 512? or the answer choices

OpenStudy (jocelm):

It's the same thing except in my mind it's a little confusing to multiply by a fraction.

OpenStudy (danjs):

you want to isolate m one one side of the equation m=.. 54 * m = 512 divide everyting by 54 54/54 * m = 512/54 1 * m = 512/54

OpenStudy (danjs):

check the numbers though that you put in the question, 512/54 is not a whole number liek the answers

OpenStudy (zeidyt):